OpenStudy (anonymous):

medal and fan! help Solve the equation for y 3y - 6x = -3

OpenStudy (anonymous):

3y - 6x = -3 y - 6x = -1 y - x = 6

OpenStudy (anonymous):

Does this help a little

OpenStudy (idealist10):

3y-6x=-3 divide by 3 y-2x=-1 y=-1+2x y=2x-1 Answer: y=2x-1